Kropf Inc. has provided the following data concerning one of the products in its standard cost system. Variable manufacturingRequired: a. Compute the materials price variance for September. b. Compute the materials quantity variance for September. C.
Answer #1

>> Standard Price = $ 7.7

>> Standard Quantity = 10,300 * 8.1 = 83,430

>> Actual Price = ( $ 667,500 / 84,700 ) = $ 7.88

>> Actual Quantity = 83,460.

_______________________________________________________________________

A. Material Price Variance = ( Standard Price - Actual Price ) * Actual Quantity

>> Material Price Variance = ( $ 7.7 - $ 7.88 ) * 83,460

>> Material Price Variance = $ 15,023 Unfavorable.

___________________________________________________________________________________

B. Material Quantity Variance = ( Standard Quantity - Actual Quantity ) * Standard Price

>> Material Quantity Variance = ( 83,430- 83,460) * $ 7.7

>> Material Quantity Variance = $ 231 Unfavorable.

================================================================================

>> Standard Rate = $ 28.7

>> Standard Hours = 10,300 * 0.4 = 4,120 Hours.

>> Actual Rate = ( $ 115,302 / 3,840 ) = $ 30

>> Actual hours = 3,840.

-----------------------------------------------------------------------------------------------------------------------------------------------------

C. Labor Rate Variance = ( Standard Rate - Actual Rate ) * Actual Hours

>> Labor Rate Variance = ( $ 28.7- $ 30) * 3,840

>> Labor Rate Variance = $ 4,992 Unfavorable.

__________________________________________________________________________________________

D. Labor Efficiency Variance = ( Standard Hours - Actual Hours ) * Standard Rate

>> Labor Efficiency Variance = ( 4,120 - 3,840) * $ 28.7

>> Labor Efficiency Variance = $ 8,036 Favorable.

======================================================================================